Roland, Mark and Peter are 3 children in the same family. Peter is 3 times Marks age. In 2 years Peter will be twice the age Roland is at that time. Roland is 1 year older than Mark. How old are the children?
Answer by ankor@dixie-net.com(22740)

Roland, Mark and Peter are 3 children in the same family.
Write an equation for each statement:
:
Peter is 3 times Marks age.
p = 3m
:
In 2 years Peter will be twice the age Roland is at that time.
p + 2 = 2(r+2)
p + 2 = 2r + 4
p = 2r + 4 - 2
p = 2r + 2
:
Roland is 1 year older than Mark.
r = m + 1
:
How old are the children?
:
In the second equation, replace p with 3m and r with m+1
p = 2r + 2
3m = 2(m+1) + 2
2m = 2m + 2 + 2
3m - 2m = 4
m = = 4 yrs is Marks age
then
r = m + 1
r = 4 + 1
r = 5 yrs is Roland's age
:
and
p = 3m
p = 3(4)
p = 12 yrs is Peter's age
:
:
Check solution in the 2nd equation
12 + 2 = 2(5+2)